the sum of two numbers is 25 and their difference is 15. what are the two numbers? let x =one number
and y= the other number. the corresponding modeling system is
{x+y=25
x-y=15
this is a system of linear equations: solution by substitution

Answer by Alan3354(69443)

x+y=25
x-y=15
================== Add
If you have to use subsititution:
x+y=25
x-y=15 --> x = y+15
Sub for x in the 1st eqn.